zarkam21:

Evaulate f(x)= 2x^2+3x+5 at f(x^2+x)

zarkam21:

This is what I did:

zarkam21:

(2x+x)(2x+x) 4x^2+2x^2+2x^2+9x^2+3x^2+3x^2+x^2

zarkam21:

:S

Vocaloid:

you're on the right track, just a few algebraic mistakes

Vocaloid:

2x^2 plugging in "(x^2+x)" for x gives us 2(x^2 + x)^2 = 2(x^2+x)(x^2+x)

Vocaloid:

let's leave the 2 in front alone for now. (x^2+x)(x^2+x) F x^2 * x^2 O x^2 * x I x * x^2 L x * x

Vocaloid:

try adding those up and combining like terms

zarkam21:

x^4+x^2+x^2+x?

Vocaloid:

be careful with exponents x^2 * x = x^3 not x^2 x * x = x^2

zarkam21:

x^4+x^2+x^3+x^2

Vocaloid:

very close the outer and inner products are the same so x^4+x^3+x^3+x^2

Vocaloid:

or, to simplify, x^4+2x^3+x^2

Vocaloid:

anyway that's the hardest part, it's easy from there on out the 3x term becomes 3(x^2+x) which is just 3x^2 + 3x then tack on the + 5 at the very end

Vocaloid:

ah, almost forgot, we had a 2 at the beginning of the binomial expansion so x^4+2x^3+x^2 becomes 2x^4 + 4x^3 + 2x^2

Vocaloid:

giving us 2x^4 + 4x^3 + 2x^2 + 3x^2 + 3x + 5 you can combine the x^2 terms as well